The volume of a cube is given by V=s^3 , where s is the length of a side. Find the length of a side of a cube if the volume is 500cm^3. Round the answer to three decimal places.
Answer by checkley77(12844)

V=s^3
500=s^3
s^3=500
s=cubert(500)
s=7.937 ans.